# HERA: High-efficiency Matrix Compression via Element Replacement

### v1

@@ Abstract
Large Language Models (LLMs) have significantly advanced natural language processing tasks such as machine translation, text generation, and sentiment analysis. However, their large size, often consisting of billions of parameters, poses challenges for storage, computation, and deployment, particularly in resource-constrained environments like mobile devices and edge computing platforms. Additionally, the key-value (k-v) cache used to speed up query processing requires substantial memory and storage, exacerbating these challenges. Vector databases have emerged as a crucial technology to efficiently manage and retrieve the high-dimensional vectors produced by LLMs, facilitating faster data access and reducing computational demands.
Effective compression and quantization techniques are essential to address these challenges, as they reduce the memory footprint and computational requirements without significantly compromising performance. Traditional methods that uniformly map parameters to compressed spaces often fail to account for the uneven distribution of parameters, leading to considerable accuracy loss. Therefore, innovative approaches are needed to achieve better compression ratios while preserving model performance.
In this work, we propose HERA, a novel algorithm that employs heuristic Element Replacement for compressing matrix. HERA systematically replaces elements within the model using heuristic methods, which simplifies the structure of the model and makes subsequent compression more effective. By hierarchically segmenting, compressing, and reorganizing the matrix dataset, our method can effectively reduce the quantization error to 12.3% of the original at the same compression ratio.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized the field of natural language processing (NLP), enabling significant advancements in tasks such as machine translation, text generation, and sentiment analysis. These models, characterized by their large-scale neural network architectures and vast training datasets, have shown remarkable capabilities in understanding and generating human language. The advent of LLMs, such as OpenAI’s GPT-3 and BERT by Google, has pushed the boundaries of what machines can achieve in linguistic tasks, providing near-human performance in various applications  Brown et al. (2020); Devlin et al. (2018).
The development of LLMs is rooted in the transformer architecture, which employs self-attention mechanisms to process and produce language with high contextual relevance. This architecture has replaced previous recurrent neural network (RNN) and long short-term memory (LSTM) models due to its efficiency and ability to handle long-range dependencies in text Vaswani et al. (2017). As a result, LLMs have become the cornerstone of modern NLP, driving innovations in areas such as automated customer service, content creation, and real-time language translation.
Furthermore, the development and deployment of LLMs face significant challenges, particularly in optimizing vector database efficiency, LLM weight quantization, and key-value (k-v) cache optimization. First, as the size and complexity of LLMs increase, efficiently managing and retrieving vectors in large databases becomes critical. Optimizing vector database efficiency involves improving indexing and search algorithms to handle high-dimensional data, which is essential for tasks like semantic search and recommendation systems Johnson et al. (2019). Second, LLM weight quantization, which involves reducing the precision of model weights, is a crucial technique for making these models more storage and computation efficient without significantly degrading their performance. Quantization can drastically reduce the memory footprint and computational requirements, enabling the deployment of LLMs on resource-constrained devices. However, achieving optimal quantization while maintaining model accuracy remains a complex challenge that requires sophisticated algorithms and techniques Dettmers et al. (2021). Lastly, optimizing k-v cache efficiency is vital for speeding up inference times in LLMs. The k-v cache stores intermediate activations during the forward pass, which can be reused to avoid redundant computations. Efficient management and compression of these caches are essential to reduce latency and improve the throughput of LLMs, especially in real-time applications like chatbots and virtual assistants.
Figure 1: The distribution of LLM parameter.
To address the challenges of storage and computational efficiency in large language models (LLMs), it is crucial to develop techniques that optimize vector database efficiency, weight quantization, and key-value (k-v) cache optimization. Therefore, finding quantization algorithms that can efficiently compress and quickly decompress matrix data is essential.
In practice, when compressing parameters, the distribution of parameters is often uneven. First, the distribution is not uniform, and second, the magnitudes of different data points vary significantly. For example, the distribution of Transformer neural network weights is shown in Figure  1. In Figure  1, different rows represent the 0th layer, 8th layer, 16th layer, and 24th layer. Different columns represent the attention layer’s q, k, v, and o layers, as well as the fully connected layer’s gate, up, and down layers. The distributions of different layers are not uniformly analyzed, and the distributions vary across different layers. Previous algorithms uniformly and linearly map parameters to another compressed space. Although this method is simple, it may not be effective and can result in significant accuracy loss because it does not take into account the original data distribution. It maps both densely populated and sparse regions to the quantized space in the same way. Besides, A lines of works have been proposed to optimize vector database efficiency Ge et al. (2013); Jegou et al. (2010); Kalantidis, Avrithis (2014), quantize the model weights Dettmers et al. (2022); Polino et al. (2018); Chmiel et al. (2020); Fan et al. (2020); Zafrir et al. (2019); Wu et al. (2022); LeCun et al. (1989) and optimizate kv cache Hooper et al. (2024); Cho et al. (2024). They also does not take into account the original data distribution. Therefore, a more reasonable approach is to quantize the model based on the distribution and magnitude of the parameters, stratifying and computing accordingly.
In this work, we propose a novel algorithm named HERA to enhance the compression of large language models. The core idea of HERA is to group the relevant matrices and compress each group using clustering methods. Furthermore, HERA reorders the original dataset based on the distribution and magnitude of the parameters. This reordering is performed in multiple layers, with each layer featuring a new arrangement that optimizes the dataset’s distribution for compression. By leveraging these techniques, HERA effectively reduces data error and minimizes the discrepancy between estimated data and actual data.
We implemented a prototype system using Python. Experimental results show that, by hierarchically segmenting, compressing, and reorganizing the matrix dataset, our method can effectively reduce the quantization error to 12.3% of the original at the same compression ratio.
2 HERA
@@ 2.1 Basic algorithm
We borrow the idea from product quantization algorithm and k-means algorithm to compress the model parameter. The algorithm is divided into four steps, the first three steps is quantization process as shown in Figure  2, the last two steps is the dequantization process as show in Figure  3. The detailed process is described below.
Algorithm 1 Product Quantization Algorithm
1:$X\in\mathbb{R}^{N\times D}$triangleright Input matrix
2:Quantization Process:
3:Step 1: Space Partitioning
4:for each vector $x_{i}$ in $X$ do
5:     Divide $x_{i}$ into $M$ low-dimensional subspaces $x_{i}^{1},x_{i}^{2},\ldots,x_{i}^{M}$
6:end for
7:Step 2: Codebook Calculation
8:for each subspace $j$ from 1 to $M$ do
9:     Apply k-means algorithm to subspace $j$ to get $k$ cluster centers
10:     Create codebook $C_{j}$ for subspace $j$ with cluster centers
11:end for
12:Step 3: Quantization
13:for each vector $x_{i}$ in $X$ do
14:     for each subspace $j$ from 1 to $M$ do
15:         Map $x_{i}^{j}$ to the nearest cluster center in $C_{j}$
16:     end for
17:end for
18:Dequantization Process:
19:Step 4: Codebook Reconstruction
20:Construct a global codebook $C$ that contains all possible combinations of cluster centers from $C_{1},C_{2},\ldots,C_{M}$
21:Step 5: Dequantization
22:for each quantized vector $q_{i}$ in the quantized dataset do
23:     Restore $q_{i}$ to its original high-dimensional space using the global codebook $C$
24:end for
  1. 1.
    Space Partitioning: Divide the high-dimensional space into a series of low-dimensional subspaces.
  2. 2.
    Codebook calculation: Use k-means algorithm to calculate codebook of low-dimensional subspaces.
  3. 3.
    Quantization: Perform independent quantization within each low-dimensional subspace, mapping data points to the nearest cluster center.
  4. 4.
    Codebook Reconstruction: Reconstruct the data using a codebook that contains all possible cluster centers.
  5. 5.
    Dequantization: Restore the vectors in the original high-dimensional space based on the vectors in the quantized low-dimensional subspaces.
Figure 2: The Quantization Process of HERA.
Figure 3: The Dequantization Process of HERA.
In data matrices where the distribution of data is uneven across different columns, directly applying the k-means algorithm for clustering may yield suboptimal results. This unevenness can distort the clustering process, as k-means assumes that all features contribute equally to the distance calculations. The differing distributions across columns mean that some features may disproportionately influence the clustering outcome, leading to biased clusters that do not accurately reflect the true structure of the data.
To mitigate these effects, it is crucial to preprocess the data appropriately. One common approach is to normalize the data so that each column contributes equally to the distance metric used by k-means. Standard normalization techniques, such as z-score normalization or min-max scaling, can be employed to transform the data into a more uniform distribution. By ensuring that all features are on a comparable scale, the k-means algorithm can perform more effectively, producing clusters that better represent the underlying patterns in the data.
Algorithm 2 Product Quantization with HERA heuristics
@@
1:Input matrix: $X\in\mathbb{R}^{N\times D}$, number of iteration in HERA: $iter$.
2:Quantized feature matrix and feature map
3:for each element $x_{i}\in X$ do
4:     Rearrange $X$ into pairs: $X_{\text{pairs}}\in\mathbb{R}^{N\times D/2}$
5:     for each dimension $j$ do
6:         $x_{pair_{i,j}}=(x_{2i,j},x_{2i+1,j})\in X_{\text{pairs}}$

### v2

@@ Abstract
The matrix quantization entails representing matrix elements in a more space-efficient form to reduce storage usage, with dequantization restoring the original matrix for use. We formulate the Quantization Error Minimization (QEM) problem as minimizing the distance between a matrix before and after quantization, under the condition that the quantized matrix occupies the same memory space. Matrix quantization is crucial in various applications, including Large Language Models (LLMs) weight quantization, vector databases, KV cache quantization, graph compression, and image compression. Recent advancements in LLMs, such as GPT-4 and BERT, have highlighted the importance of matrix compression due to the large size of parameters and KV cache, which are stored as matrices.
We propose HETA (HETA) to address the QEM problem by leveraging the local orderliness of matrix elements, involving iterative element swapping to form a locally ordered matrix. This matrix is then grouped and quantized by columns. To enhance HETA, we introduce two optimizations: further quantizing residuals to reduce MSE, and using masking and batch processing to accelerate the algorithm.
Experimental results demonstrate that HETA can effectively reduce MSE to 12.3% of the original value at the same compression ratio, outperforming the best baseline algorithms. Our contributions include the abstraction of the QEM problem, the design of the HETA algorithm, and the proposal of two optimizations to improve accuracy and speed.
@@ 1 Introduction
The matrix quantization entails representing the elements of a matrix in a more space-efficient form to reduce its storage usage. Dequantization, on the other hand, is the process during usage where the original matrix is restored from the quantized matrix using a restoration algorithm. We formulate the Quantization Error Minimization (QEM) problem as the task of minimizing the distance between a matrix before and after quantization in high-dimensional space, under the condition that the quantized matrix occupies the same space.
Matrix quantization is widely employed across diverse applications, including Large Language Models (LLMs) weight quantization (Dettmers et al. (2024); Lin et al. (2024); Shao et al. (2023); Xiao et al. (2023)), vector database (Xu et al. (2018); Jegou et al. (2010)), LLM k-v cache quantization (Liu et al. (2024); Zhang et al. (2024); Hooper et al. (2024); Kawakibi Zuhri et al. (2024); Duanmu et al. (2024); Yue et al. (2024); Lee et al. (2024); Adnan et al. (2024)), graph compression (Brisaboa et al. (2009); Claude & Ladra (2011)), and image compression (Yu et al. (2018); Ning et al. (2016)).
@@
Specifically, the recent advancements in Large Language Models (LLMs) have made matrix compression even more critical. Large Language Models (LLMs) have revolutionized the field of natural language processing (NLP), enabling significant advancements in tasks such as machine translation, text generation, and sentiment analysis. These models, characterized by their large-scale neural network architectures and vast training datasets, have shown remarkable capabilities in understanding and generating human language. The advent of LLMs, such as OpenAI’s GPT-4 (Achiam et al. (2023)), and BERT (Devlin et al. (2018)) by Google, has pushed the boundaries of what machines can achieve in linguistic tasks, providing near-human performance in various applications. The parameters and KV cache in Large Language Models are very large in size. For instance, GPT-2 contains 1.5 billion parameters (Solaiman et al. (2019)), whereas GPT-3 has expanded to 175 billion parameters (Brown et al. (2020)). Additionally, the KV cache accounts for over 30% of GPU memory during deployment, compared to the 65% occupied by the parameters (Kwon et al. (2023)). Both the parameters and KV cache are stored in the form of matrices. GPTQ (Frantar et al. (2022)) and SqueezeLLM (Kim et al. (2023)) directly address the QEM problem by treating it as the optimization objective for quantizing parameters and the KV cache. Therefore, the abstracted scientific problem of QEM is highly significant.
There has been a growing literature on the matrix quantization. The first category of methods focuses on independently compressing the elements of a matrix for each specific scenario. Examples of such techniques include LLM.int8 (Ge et al. (2013)), Optimal Brain Damage (LeCun et al. (1989)), GPTQ (Frantar et al. (2022)), AWQ (Lin et al. (2023)), SmoothQuant (Xiao et al. (2023)), and OmniQuant (Shao et al. (2023)). The second category groups matrix elements by columns and then applies quantization to each group. Relevant works in this category include product quantization (PQ) (Jegou et al. (2010)), optimized product quantization (OPQ) (Ge et al. (2013)), locally optimized product quantization (LOPQ) (Kalantidis & Avrithis (2014)), and SqueezeLLM (Kim et al. (2023)). The first category of work can be summarized as utilizing outliers and the importance of elements in specific scenarios to improve the RTN (Round-To-Nearest)  [^1] RTN (Round-To-Nearest) is an algorithm that quantizes elements by rounding each value to its nearest representable level. (Gray & Neuhoff (1998) algorithm). The second category focuses on enhancements to PQ algorithms. The primary drawback of the first category is that it does not consider the correlation between elements, instead independently quantizing and storing elements. Conversely, the second category fails to account for the relative order of elements.
In our research, we observed that the order of elements has a significant impact on the quantization outcome. Intuitively, consider a matrix $\begin{bmatrix}1&2\\ 2&1\end{bmatrix}$. To losslessly quantize this matrix (with MSE=0), two vectors, $[1,2]$ and $[2,1]$, are required. However, if the matrix is reordered to $\begin{bmatrix}1&2\\ 1&2\end{bmatrix}$, it can be losslessly quantized using just a single vector, $[1,2]$.
Using these ideas, we propose HETA (HETA) to address the QEM problem in matrix quantization. The core idea of HETA is to leverage the local orderliness of matrix elements to optimize the QEM problem. The design involves swapping adjacent elements of the matrix to form a new, locally ordered matrix. To cover a broader range, we can perform further element swapping on the initially locally ordered matrix, similar to the approach of receptive field used in convolutional neural networks (LeCun et al. (1998)). This step can be repeated for multiple iterations. The newly ordered matrix is then grouped by columns, with each group being quantized.
Additionally, we propose two optimization algorithms based on the basic HETA. First, the residuals between the original and quantized matrices can be further quantized to enhance the accuracy of the results. Second, since the iterative swapping operations can slow down the algorithm, we introduce optimization techniques based on masking and batch processing to accelerate the HETA.
We evaluate HETA on multiple datasets. Experimental results demonstrate that our method can effectively reduce the MSE to 12.3% of the original value at the same compression ratio, compared to the best baseline algorithm, which achieves an MSE of XXXX%.
We summarize our contributions below.
  • Abstracted a problem: We abstracted the Quantization Error Minimization (QEM) problems from various application scenarios.
  • Designed an algorithm: We developed the HETA (HETA) algorithm, leveraging the concept of local orderliness to optimize the QEM problem.
  • Proposed two optimizations: We introduced the use of residuals to reduce MSE and employed batch processing techniques to accelerate HETA algorithm.
2 PROBLEM SETTING
In this section, we formally define the Quantization Error Minimization (QEM) problem and analyze the impact and feasibility of rearranging matrix elements prior to quantization.
QEM Problem: we denote the matrix to be quantized as $X$, the quantized matrix as $X^{q}$, and the matrix obtained by dequantizing $X^{q}$ as $X^{\prime}$. Both $X$ and $X^{\prime}$ are $n\times d$ dimensional, but the dimensions of $X^{q}$ depend on the quantization algorithm. The elements of these matrices are denoted as $x_{(i,j)}$, $x^{q}_{(i,j)}$, and $x^{\prime}_{(i,j)}$, respectively. The memory occupied by a matrix is denoted as memory(), and the memory constraint as mem_constrain.
Definition 1 (Quantization Error Minimization (QEM) Problem).
The Quantization Error Minimization (QEM) problem is defined as the task of minimizing the combined objective of the Mean Squared Error (MSE) and the memory size of the quantized matrix:
$\text{minimize}\;\text{MSE}(X,X^{\prime})+\lambda\cdot\text{memory}(X^{q}),$
where
$\text{MSE}(X,X^{\prime})=\frac{1}{n\cdot d}\sum_{i,j}\left(x_{(i,j)}-x^{\prime}_{(i,j)}\right)^{2},$
$\lambda$ is a regularization parameter that balances the trade-off between minimizing the MSE and the memory usage of the quantized matrix.
Subject to the condition that the quantized matrix occupies the same space:
[…]
Rearranging matrix elements: Our idea is to rearrange the matrix elements before quantization. Let the quantization algorithm be denoted as quant() and the rearrangement algorithm as rearrange(). Previously, the quantized matrix was obtained as $X^{q}=\text{quant}(X)$, whereas in our rearrangement approach, it is obtained as $X^{q}=\text{quant}(\text{rearrange}(X))$. Therefore, the rearrangement algorithm only affects the quantization if the quantization method is order-sensitive.For instance, order-insensitive algorithms like RTN do not benefit from rearrangement because RTN processes each element independently, while order-sensitive algorithms like PQ can be optimized through rearrangement.
However, finding the optimal arrangement requires exploring a large search space. Specifically, for an $n\times d$ matrix, there are $(n!)^{d}$ possible rearrangements. Therefore, searching the entire space is impractical, and we need heuristic algorithms to prioritize exploring more promising regions of the search space.
3 METHODS
The core idea of HETA (HETA) is to leverage the local orderliness of matrix elements by rearranging adjacent elements to optimize the matrix quantization algorithm. Additionally, multiple iterations are used to expand the algorithm’s swapping field. We begin by describing the basic version of this idea and then propose two optimizations for the basic algorithm.
@@ 3.1 Basic algorithm
The steps and design of the HETA algorithm are illustrated in Figure 1. The figure provides a conceptual overview of the HETA (HETA) algorithm.
On the left side of the figure, the process of rearranging matrix elements through multiple iterations in the HETA is depicted. This iterative procedure involves comparing adjacent elements and ensuring that adjacent elements are not placed in the same matrix. This separation of adjacent elements into different matrices is conceptually similar to quantum entanglement, hence the name ”HETA.” The HETA algorithm performs multiple iterations, with each iteration comparing adjacent elements and splitting them based on their values. This approach is based on two observations. First, the local orderliness of the matrix can enhance the regularity of matrix elements, thereby improving compression efficiency. Second, multiple iterations allow the orderliness to cover a larger number of elements. For instance, adjacent elements after the first iteration might be non-adjacent in the original matrix. This iterative process, similar to that of convolutional neural networks, extends the ”receptive field” to cover more elements.
On the right side of the figure, the HETA compression process is shown. After the elements have been rearranged, the algorithm groups the sub-matrices, followed by clustering to determine the centroids. These centroids are used to quantize the matrix. The final quantized matrix, codebook, and indicator maps are the outputs of the HETA algorithm. This divide-and-conquer compression method is based on an observation: by splitting the matrix, the expressive power can be increased under a fixed storage constraint. For example, if the matrix is divided into $m$ blocks, each with $c$ centroids, the entire matrix can express $c^{m}$ centroids while only storing $c\times m$ centroids. Without splitting, storing $c^{m}$ centroids would be required.
3.2 HETA Quantization
The proposed HETA Algorithm is designed to optimize the quantization of a matrix $X\in\mathbb{R}^{n\times d}$ by leveraging local orderliness. The algorithm proceeds in four key steps, as shown in Algorithm 2:
Step 1: Initial Partitioning. The algorithm starts by comparing each adjacent pair of elements $(x_{i,j},x_{i,j+1})$ in the input matrix $X$. Based on the comparison of these elements, the smaller value is placed into matrix $S$ and the larger into matrix $L$. Concurrently, an Indicator Map $I$ is generated, where a 0 denotes that the smaller element is on the left and a 1 denotes it is on the right. This process results in the creation of two matrices, $S$ and $L$, which collectively represent the initial partitioning of the original matrix $X$.
Step 2: Recursive Partitioning. Following the initial partitioning, the algorithm undertakes a recursive partitioning process. Starting from $k=1$, the matrices $S^{i}_{k}$ and $L^{i}_{k}$ [^2] Since each level has multiple $S$ and $L$ matrices, we use the subscript to denote the iteration number and the superscript to denote the sequence within each iteration. derived from the previous iteration (for $i=1,2,\dots,2^{k-1}$) are further partitioned based on their sizes. In each iteration, a new Indicator Map $I_{k+1}$ is generated and then merged with the existing Indicator Maps. This recursive process continues until the predefined number of iterations $l$ is reached. The resulting Indicator Maps $I_{1},I_{2},\dots,I_{l}$ are stored in the set $IM$. The final matrices $S_{l}$ and $L_{l}$ are combined to form a new matrix $X^{*}$, which represents a locally ordered version of the original matrix.
Step 3: Subspace Grouping. After constructing the matrix $X^{*}$, it is divided into subspaces $\{G_{i}\}$, where $i=1,2,\dots,m$. This division is based on the local ordering of elements, which facilitates more effective clustering and quantization in the subsequent steps.
Step 4: Clustering and Quantization. In the final step, each subspace $G_{k}$ (for $k=1,2,\dots,m$) is clustered to determine a set of centroids $C_{k}=\{\mathbf{c}_{j}\}$. Each vector $\mathbf{g_{i}}$ within the subspace is then quantized by assigning it to the nearest centroid. This process results in the quantized matrix $X^{q}$. The output of the algorithm includes the quantized matrix $X^{q}$, the codebook $\mathbf{C}$ containing the centroids, and the Indicator Maps $IM$.
Algorithm 1 HETA Quantization Algorithm
@@ 3.2 HETA Quantization
1:Input: Matrix $X\in\mathbb{R}^{n\times d}$, $m$ is the number of subspaces for quantization, $l$ is the number of HETA iterations.
2:Output: Quantized matrix $X^{q}=\{X^{q}_{i,k}\}$ where $i=1,\dots,d$ and $k=1,\dots,m$, Codebook $\mathbf{C}=\{C_{i}\}$ where $i=1,\dots,m$, and Indicator Maps $IM=\{I_{i}\}$ where $i=1,\dots,l$
3:Initialize $IM$ and $X^{q}$ as empty
4:Step 1: Initial Partitioning
5:for each adjacent pair $(x_{i,j},x_{i,j+1})$ in $X$ do
6:     if $x_{i,j}>x_{i,j+1}$ then
7:         $S_{i,j}\leftarrow x_{i,j+1}$, $L_{i,j}\leftarrow x_{i,j}$
8:         $I\left(\frac{i}{2},j\right)\leftarrow 0$
9:     else
10:         $S_{i,j}\leftarrow x_{i,j}$, $L_{i,j}\leftarrow x_{i,j+1}$
11:         $I\left(\frac{i}{2},j\right)\leftarrow 1$
12:     end if
13:end for
14:Step 2: Recursive Partitioning
15:$k\leftarrow 1$
16:while $k\neq l$ do
17:     for $i=1$ to $2^{k-1}$ do
18:         Apply size-based partitioning to $S^{i}_{k}$ and $L^{i}_{k}$
19:         Generate and merge new Indicator Maps into $I_{k+1}$ for each partition
20:     end for
21:     $k\leftarrow k+1$
22:end while
23:Store $I_{1},I_{2},\dots,I_{l}$ into $IM$
24:$X^{*}\leftarrow S^{1}_{l}\cup L^{1}_{l}\cup S^{2}_{l}\cup L^{2}_{l}\cup\dots\cup S^{2^{l-1}}_{l}\cup L^{2^{l-1}}_{l}$triangleright Combine the locally ordered parts from the final layer $l$
25:Step 3: Subspace Grouping
26:Group the matrix $X^{*}$ into subspaces $\{G_{i}\}$ where $i=1,2,\dots,m$
27:Step 4: Clustering and Quantization
28:for each vector $\mathbf{g_{i}}$ in $G_{k}$ (for $k=1,2,\dots,m$do
29:     Apply clustering to find centroids $C_{k}=\{\mathbf{c}_{j}\}$
30:     $X^{q}_{i,k}\leftarrow\arg\min_{\mathbf{c}_{j}}\|\mathbf{g_{i}}-\mathbf{c}_{j}\|$triangleright Quantize to nearest centroid within group
31:end for
32:Return Quantized matrix $X^{q}$, Codebook $\mathbf{C}$, and Indicator Maps $IM$
3.2.1 HETA Dequantization
The dequantization process, as described by the HETA Dequantization Algorithm, aims to reconstruct the original matrix $X^{\prime}$ from its quantized version $X^{q}$. This process involves several key steps, as shown in Algorithm 2:
Step 1: Dequantization Using Codebook. The first step involves using the codebook $\mathbf{C}$ to directly retrieve the original vectors corresponding to each quantized vector $\mathbf{g_{i}^{q}}$ in the matrix $X^{q}$. For each vector $\mathbf{g_{i}^{q}}$ in $X^{q}_{i,k}$, the algorithm assigns $\mathbf{g_{i}^{\prime}}$ to the centroid $\mathbf{c}_{j}$ from the codebook, where $j$ is the index corresponding to the quantized vector. This step effectively reverses the quantization by mapping the quantized vectors back to their original forms using the codebook entries.
Step 2: Reverse Subspace Grouping. After mapping, the algorithm combines the dequantized subspaces $\{G_{i}^{\prime}\}$, where $i=1,2,\dots,m$, into a single matrix $X^{\prime*}$. Each subspace $G_{i}^{\prime}$ contains the vectors $\mathbf{g_{i}^{\prime}}$ corresponding to the original matrix entries. This step reconstructs the subspaces into an uniform matrix that is a preliminary version of the original matrix $X^{\prime}$.
Step 3: Reverse Recursive Partitioning and Final Reconstruction. The final step involves reversing the recursive partitioning process that was applied during quantization. Starting from the last iteration $k=l$, the algorithm merges the matrices $S_{k}$ and $L_{k}$ based on the Indicator Maps $I_{k}$ and the matrix $X^{\prime*}$. This merging process is iteratively applied from $k=l$ down to $k=1$, updating $X^{\prime*}$ at each step to gradually reconstruct the original matrix structure. The final output of the algorithm is the reconstructed matrix $X^{\prime}\in\mathbb{R}^{n\times d}$, which approximates the original matrix before quantization.
Algorithm 2 HETA Dequantization Algorithm
1:Input: Quantized matrix $X^{q}=\{X^{q}_{i,k}\}$, where $i=1,\dots,d$ and $k=1,\dots,m$; Codebook $\mathbf{C}=\{C_{i}\}$, where $i=1,\dots,m$; and Indicator Maps $IM=\{I_{i}\}$, where $i=1,\dots,l$
2:Output: Reconstructed matrix $X^{\prime}\in\mathbb{R}^{n\times d}$
3:Initialize $X^{\prime}$ as empty
4:Step 1: Dequantization Using Codebook
5:for each vector $\mathbf{g_{i}^{q}}$ in $X^{q}_{i,k}$ do
6:     $\mathbf{g_{i}^{\prime}}\leftarrow\mathbf{c}_{j}$, where $j$ corresponds to the codebook entry for $\mathbf{g_{i}^{q}}$triangleright Direct lookup from codebook